a worker drops a wrench from the top of a tower 80 meters tall. What is the velocity when the wrench strikes the ground assuming the acceleration to be -9.81 m/s^2
1
Expert's answer
2012-09-07T10:34:26-0400
At first, let's find the time needed to fall down from the height of 80 meters
from the equation
for displacement (assume the axis is directed downwards and
the origin is on the top of a tower):

h = g * t^2 / 2,
t = sqrt(2h /
g).

Now let's use the equation for the speed when the wrench falls
down:

V = g * t = g * sqrt(2h / g) = sqrt(2 * g * h).
So,
V =
sqrt(2 * 9.81 * 80) = 39.6 m/s.